Atoms of CO2 gas decay from upper to lower level with a spread in frequency is about
10kHz. Calculate the line shape if the broadening mechanism is found Doppler
broadening ?
1
Expert's answer
2020-05-21T13:35:02-0400

The line shape is Gaussian. The distribution of the frequencies of the absorbed radiation is


"L=\frac{1}{\sigma\sqrt{2\pi}}\text{exp}\bigg(-\frac{(f-f_0)^2}{2\sigma^2}\bigg),\\
\sigma=\frac{f_0}{c}\sqrt{\frac{2kT\text{ ln}2}{m}}."

In the equation above m is the mass of molecules and T - absolute temperature of gas.